Predict the mass of hydrogen gas produced when 2.73 g of aluminum reacts in a single replacement reaction with excess sulphuric acid.
1
Expert's answer
2013-01-16T12:11:00-0500
2Al+ 3H2SO4 = Al2(SO4)3 + 3H2

First, look for the amount of substance aluminum oxide:
n(Al2O3)=m/M
n(Al2O3)= 100g/ 102g/mol = 0.9999 mol
1 mol Al2O3 ----> 2mol Al
n(Al)= 1,9998 mol
m(Al)= n*M
m(Al) = 1,9998mol*27 g/mol = 53,99 g

Answer:
53,99g